Information from the financial statements of Ames Fabricators, Incorporated included the following:


December 31
20242023
Common shares100,000100,000
Convertible preferred shares (convertible into 68,000 shares of common)26,40026,400
8% convertible bonds (convertible into 30,000 shares of common)$ 1,000,000$ 1,000,000

Ames’s net income for the year ended December 31, 2024, is $740,000. The income tax rate is 25%. Ames paid dividends of $5 per share on its preferred stock during 2024.

Required:

Compute basic and diluted earnings per share for the year ended December 31, 2024.

Note: Do not round intermediate calculations. Enter your answers in thousands (i.e., 10,000 should be entered as 10).
Answer #1

Answer:


numerator/denominator=earnings per share
basic608/100=6.08
diluted800/198=4.04

explaination:

step1) 740,000-(5*26,400)=608,000

step 2) 608,000/100= 6.08


step3) convertible preferred stock 5*26,400= 132,000

convertible bonds (1,000,000*8%)*(1-25%)= 60,000

step 4) 608,000+132,000= 740,000+60,000 shares= 800,000

100,000+68,000= 168,000+ 30,000 shares= 198,000

800,000/198,000= 4.0404040404 or 4.04


answered by: anonymous